OpenStudy (anonymous):

square root of 72x^3y^4z^9

OpenStudy (saifoo.khan):

6root2 x^3/2 root3y^2 2z^3

OpenStudy (anonymous):

I understand how you got 6root2x^3/2 but I don't understand root 3y^2 2z^3

OpenStudy (saifoo.khan):

http://www.twiddla.com/562069

OpenStudy (saifoo.khan):

come here

jhonyy9 (jhonyy9):

(72x*3y*4z*9)*(1/2) =(36)(2)x*2(x)(y*2)*2(z*4)*2(z) =6xy*(2)z*(4)radical(2xz)